17. If the confidence level for a confidence interval is 98%. then the error rate α is ______. a. 98% b. 2% c. 95% d. 5% 18. Suppose the sample mean is μ and the confidence interval for population mean is (μ-E, μ+E). E will be getting smaller if: (18.a) the sample size is getting ___________; a. smaller b. larger (18.b) the standard deviation of the data in the sample is __________; a. smaller b. larger (18.c) the allowed error rate α is ___________; a. smaller b. larger (18.d) the confidence level (1-α) is ___________. a. lower b. higher.
Answer #1

17)then the error rate α is =1 - 0.98 = 0.02= 2% ( option b)

18) E = z (s/ sqrt (n))

E will be getting smaller if

a) sample size larger

b) the standard deviation of the data in the sample is smaller

c) the allowed error rate α is larger

d)the confidence level (1-α) is lower